When the sum of square residual is 0, R2 is

Need more information to answer this question.

a. 0

b. 0.5

c. 1
Answer #1

Answer

We know that R^2 is given as

R^2 =1 - (SSE/SST)

where SSE is sum of square residual and SST is sum of squares total

when SSE = 0, we get

R^2 = 1- (0/SST)

= 1- 0

= 1

So, option C is correct

Answer is R^2 = 1